OpenStudy (anonymous):

Solve systems numerically. x^3=4x^3-2 I am stuck and have no idea what to do next. I know I have to eventually find a point (x,y) where they intersect.

OpenStudy (anonymous):

move x in one side so x^3-4x^3=-3

OpenStudy (anonymous):

just add now

OpenStudy (anonymous):

Okay, so I add 3 to what though? x^3-4x^3?

OpenStudy (anonymous):

-3x^3=-2

OpenStudy (anonymous):

divid by -3 both side

OpenStudy (anonymous):

*sides

OpenStudy (anonymous):

Okay I will, one minute

OpenStudy (anonymous):

I got x^3=2/3. Is that correct?

OpenStudy (anonymous):

yes..... now take cubic squr.

OpenStudy (anonymous):

cubic square?

OpenStudy (anonymous):

sorry cubic root

OpenStudy (anonymous):

\[x=\sqrt[3]{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

Once you have x, don’t forget to find y (or f(x) ), since you should be finding a point (x,y) where these intersect. This is what my teacher said, if that helps.

OpenStudy (anonymous):

\[x ^{3}=\frac{ 2 }{ 3 }\] taking the cubic root for both sides \[\sqrt[3]{x ^{3}}=\sqrt[3]{\frac{ 2 }{ 3 }}\] \[x=\sqrt[3]{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

but where is y in the function!

OpenStudy (anonymous):

There is none so I think she meant to just find f(x)

OpenStudy (anonymous):

The original problem was f(x)=x^3 and g(x)=4x^3-2. And it just said to solve the systems numerically. And there is supposed to be a point where (x,y) they intersect.

OpenStudy (anonymous):

okay then you don't need to solve for x since x is cubed. you just let x^3=2/3. use this value in f(x)

OpenStudy (anonymous):

Okay, so how exactly do I start solving for f(x)?

OpenStudy (anonymous):

f(x)=x^3 so if x=0, f(x)=0 x=2, f(x)=8

OpenStudy (anonymous):

Do I need to solve any further or are those just the (x,y) points?

OpenStudy (anonymous):

but using the value that you go 2/3 is where they should intersect

OpenStudy (anonymous):

no \[x=\sqrt[3]{\frac{ 2 }{ 3 }}\] f(x)=2/3 g(x)=4(2/3)-2=8/3-2=(8-6)/3=2/3

OpenStudy (anonymous):

Okay, so those are the final answers then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, thank you so much for taking the time to help me. I was so confused. You're a life saver.